# Temporal Feature Matters: A Framework for Diffusion Model Quantization

### v3

@@ 4.1 Disturbance Observations
Figure 2: FID [21] on LSUN-Bedrooms 256$\times$256 [95] for LDM-4, with Gaussian noise applied to its temporal/non-temporal features activations.
Temporal feature disturbance. Having established the sensitivity of the temporal feature to disturbance, we thoroughly analyze its variations before and after the quantization of embedding layers and time embed in the Stable Diffusion model. Prior to this analysis, we introduce the temporal/non-temporal feature activation error at $t$ as defined by:
@@
Figure 3: (Left) Temporal feature disturbance. The green inflection points highlight the significant phenomenon of temporal feature disturbance. (Right) Temporal information mismatch ($i=11$). The coordinates of the inflection points on the blue curve can denoted as $(t,t+\delta_{t,i})$. It indicates $\widehat{\mathbf{X}_{t,i}^{temporal}}$ exhibits the highest similarity with $\mathbf{X}_{t+\delta_{t,i},i}^{temporal}$. To be noted, for any point on the polylines in the above figures, assuming its x-axis is $t_{0}\in[1,\ldots,T]$, we pursue $T-t_{0}+1$ denoising processes to obtain its y-axis from the same random noise. Specifically, the model is set to full precision from $T$ to $t_{0}+1$ and only quantized precision at $t_{0}$.
@@ 4.2 Disturbance Impacts
Temporal information mismatch. Obviously, temporal feature disturbance alters the original embedded temporal information. Specifically, $\mathbf{X}_{t,i}^{temporal}$ is intended to correspond to time-step $t$. However, due to significant errors, the quantized model’s $\widehat{\mathbf{X}_{t,i}^{temporal}}$ is no longer accurately associated with $t$, resulting in what we term as temporal information mismatch:
$t\leftarrow\mathbf{X}_{t,i}^{temporal},\quad t\nleftarrow\widehat{\mathbf{X}_{t,i}^{temporal}}.\$ (9)
Furthermore, as depicted in Fig. 3 (Right), we even observe a pronounced temporal information mismatch in Stable Diffusion. Specifically, the temporal feature generated by the quantized model at time-step $t$ exhibits a divergence from that of the full-precision model at the corresponding time-step. Instead, it tends to align more closely with the temporal feature corresponding to $t+\delta_{t}$, importing wrong temporal information from $t+\delta_{t}$.
@@
where we apply disrupted temporal features to the model. Evidently, the deviation in the denoising trajectory intensifies as the number of denoising iterations increases, resulting in the final generated image failing to align accurately with $\mathbf{x}_{0}$. This evolution is illustrated in Fig. 4, where we maintain UNet in full precision, except for embedding layers and time embed.
We define temporal feature disturbance as a significant temporal feature error (Sec. 4.1). Then we show that temporal feature disturbance (i.e., significant temporal feature error) accounts for temporal information mismatch (Sec. 4.2), which leads to trajectory deviation (Sec. 4.2). In Fig. 4, the deviated trajectory can result in noisy images gradually shifting from a normal denoising trajectory. Thus, the image quality degrades severely. Moreover, as reflected in Fig. 2, with temporal feature error increasing [^4] Obviously, temporal feature error which can also be defined by $\mathcal{L}^{temporal}_{t,i}$ in Eq. (13) increases with the increase of noise level $\lambda$., generation quality measured by FID score deteriorates sharper and sharper. Sec. A for a more detailed relationship between temporal feature errors and final generation quality.
@@
Figure 4: Denoising process of full-precision (Upper) and w4a8 quantized (Lower) Stable Diffusion $(T=50)$ under the same experiment settings, $\mathbf{x}_{0}$. It is noteworthy that, in the quantized model employed here, to showcase the impact of temporal features, only the layers generating temporal features are quantized and the components unrelated to the generation of temporal features are maintained in full precision.
@@ 4.3 Inducement Analyses
TABLE I: FID, sFID and SQNR on LSUN-Bedrooms $256\times 256$ [95] for LDM-4. “Prev” represents BRECQ. Freeze denotes our trial. We conduct the same experiments with more baselines as “Prev” in Sec. B, which further validate our analysis.
@@
Unaware of finite activations within $h(\cdot)$ and $g_{i}(\cdot)$. We observe that, given $T$ as a finite positive integer, the set of all possible activation values for embedding layers and time embed is finite and strictly dependent on time-steps. Within this set, activations corresponding to the same layer display notable range variations across different time-steps  [^5] The details of range variations can be found in Sec. B in the appendix of [24]. Previous methods [81, 85] mainly focus on finding the optimal calibration method for $\widehat{\mathbf{x}}_{t}$-related network components. Moreover, akin to the first inducement, their calibration is directly towards the Residual Bottleneck Block, which proves suboptimal  [^6] The evidence can be found in Sec. C in the appendix of [24]. However, based on the finite activations, we can employ calibration methods, especially for these time information-related activations, to better adapt to their range variations.
4.4 Quantization Framework
To mitigate temporal feature disturbance, we propose two different temporal feature maintenance strategies categorized into TIB-based and cache-based ones in Sec. 4.4.1 and Sec. 4.4.2, respectively. Finally, we incorporate both scenarios together with our selection strategy to further solve the problem in Sec. 4.4.3. Our methods are outlined in Fig. 5.
4.4.1 TIB-based Maintenance
@@
$\mathop{\arg\,\min}\limits_{\mathcal{S}^{\mathbf{W}}_{g\circ h},\mathcal{Z}^{\mathbf{W}}_{g\circ h}}\sum_{i=1}^{n}\underbrace{\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2}}_{{\color[rgb]{0,0,0}\mathcal{L}^{temporal}_{t,i}}},$ (11)
@@
Finite Set Calibration. To address the challenge posed by the wide span of activations within a finite set for the second inducement, we propose Finite Set Calibration (FSC) for activation quantization. This strategy employs $T$ sets of quantization parameters for activation, such as $\{(s_{T},z_{T}),\ldots,(s_{1},z_{1})\}$ for an arbitrary activation $\mathbf{X}$ within embedding layers and time embed. At time-step $t$, the quantization function for the $\mathbf{X}$ can be expressed as:
@@
where $s_{t}$ and $z_{t}$ are the corresponding scale and zero offset. Equipped with FSC, we can obtain $\mathcal{S}^{\mathbf{X}}_{g\circ h}$ and $\mathcal{Z}^{\mathbf{X}}_{g\circ h}$, which refer to all the quantization scales and zero offsets for the activations  [^8] We consider the outputs and inputs of every linear operator like previous works [79, 19]. in $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$, respectively. To be noted, the calibration target for these activations is also aligned with the output of the TIB. Besides that, we find that Min-max [62] can achieve satisfactory results with high efficiency (more evidence in Sec. 5.4.2) for range estimation. In Sec. 5.3.1, we perform a detailed analysis to demonstrate that our method incurs a negligible extra cost for inference.
4.4.2 Cache-based Maintenance
@@
where $s_{t,i}$ and $z_{t,i}$ denote the quantization scale and zero offset for the pre-gained $\mathbf{X}_{t,i}^{temporal}$ (i.e., the output of $g_{i}\circ h$ at $t$). Finally, the obtained quantization parameters can be defined as $\mathcal{S}^{temporal}=\{s_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$ and $\mathcal{Z}^{temporal}=\{z_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$. Specifically, in this strategy, we employ LSQ [12] to optimize every objective  [^9] From Eq. (13), it is obvious that we have to optimize $n\times T$ objectives across different indexes of temporal features. separately, which can be done in just a few minutes. During inference, this strategy enables us to get the temporal feature by reloading the cached items. In addition, we find that this approach can help improve latency in cases. More detailed analysis can be found in Sec. 5.3.2.
@@
4.4.3 Disturbance-aware Selection
Figure 6: Comparison of the temporal feature disturbance between TIB-based and Cache-based Maintenance with w4a8 quantized Stable Diffusion ($T=50$).
@@ 4.4.3 Disturbance-aware Selection
Figure 7: Samples from w4a8 quantized Stable Diffusion on MS-COCO [48] caption with TIB-based Maintenance (Left), Cache-based Maintenance (Middle), and Disturbance-aware Selection (Right), where green rectangles present finer points than red rectangles. Numbers in the last row denote FID on MS-COCO caption of the corresponding methods.
@@
5 Experiments
We organize the experiments as follows: In Sec. 5.1, we first demonstrate implementation details. Then, we exhibit the outstanding performance gained by our framework in Sec. 5.2. Further, we also deploy our quantized models on various hardware to investigate their inference efficiency in Sec. 5.3. Last, we conduct comprehensive ablative studies for the proposed framework in Sec. 5.4. It is worth noting that we also discuss and conduct experiments for video generation in Sec. C. A guideline on when to use which proposed strategies and sub-4-bit quantization results are included in Sec. D and Sec. E, respectively.
5.1 Implementation details
Models and datasets. In this section, we conduct image generation experiments to evaluate the proposed quantization framework on various diffusion models: pixel-space diffusion model DDPM [22] for unconditional image generation, latent-space diffusion model LDM [74] for unconditional image generation and class-conditional image generation. We also apply our work to Stable Diffusion-v1-4 [74], SD-XL-base-1.0 [71], and SD-XL-turbo [78] for text-guided image generation. In our experiments, We use seven standard benchmarks: CIFAR-10 $32\times 32$ [41], LSUN-Bedrooms $256\times 256$ [95], LSUN-Churches $256\times 256$ [95], CelebA-HQ $256\times 256$ [30], ImageNet $256\times 256$ [10], FFHQ $256\times 256$ [32] and MS-COCO [48].
@@ 5.1 Implementation details
Evaluation metrics. For each experiment, we evaluate the performance of diffusion models with Fréchet Inception Distance (FID) [21], and Signal-to-quantization-noise Ratio (SQNR) [67] which measures the quantization distortion in detail. In the case of LDM and text-guided generation experiments, we also include sFID [76], which better captures spatial relationships than FID. For ImageNet and CIFAR-10 experiments, we additionally provide Inception Score (IS) [76] as a reference metric. Further, in the context of text-guided experiments, we extend our evaluation to include the compatibility of image-caption pairs, employing the CLIP score [20]. The ViT-B/32 [11] is used as the backbone when computing the CLIP score. To ensure consistency in the reported outcomes, all results are derived from our implementation or from other papers, where experiments are conducted under conditions consistent with ours. More specifically, in the evaluation process of each experiment, we sample 50k images from DDPM or LDM, 30k images from Stable Diffusion or SD-XL-turbo, and 10K images from SD-XL. All experiments are conducted utilizing one H800 GPU and implemented with the PyTorch framework [68] without special claims.
5.2 Performance Comparison
Unconditional generation. In the experiments conducted on the LDM with DDIM sampler [82], we maintain the same experimental settings as presented in  [74], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. II, the FID performance differences relative to the full precision (FP) model are all within 0.7 for all settings employing DS. Specifically, on the CelebA-HQ $256\times 256$ dataset, DS exhibits an FID reduction of 6.74, an sFID reduction of 6.79, and an SQNR increase of 4.09 in the w4a8 setting compared to the current algorithms. It is noticeable that existing methods, whether in 4-bit or 8-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TM/CM/DS shows almost no performance degradation. Importantly, DS also achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the w4a8 setting, with FID and sFID reductions of 2.33 and 7.67 compared to PTQD [19], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of 8 may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely than existing methods.
@@ 5.2 Performance Comparison
Besides the experiments on LDM, We have also conducted experiments with DDPM on CIFAR-10 $32\times 32$. As shown in Tab. III, our methods still achieve comprehensive improvements in terms of IS, FID, and SQNR compared to the existing advanced methods. However, due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation, so the results we obtain may not be as pronounced.
Class-conditional generation. On the ImageNet $256\times 256$ dataset, we employed a denoising process with 20 steps DDIM sampler [82], setting eta and cfg to 0.0 and 3.0, respectively. In Tab. IV, compared to PTQD, our method achieves a FID reduction of 2.09 on w8a8, and a 5.35 sFID decrease on w4a8. Under all the conditions with DS, we observe an improvement of over 9 in IS. Particularly noteworthy is that, across various quantization settings, our method consistently achieved lower FID and higher SQNR compared to the FP model.
TABLE IV: Quantization results for unconditional image generation with class-conditional image generation with LDM-4 on ImageNet $256\times 256$.
@@
TABLE V: Quantization results for text-guided image generation with Stable Diffusion-v1-4, SD-XL-base-1.0, and SD-XL-turbo on MS-COCO captions.
@@
“Elephant displayed in window of city building near roadway. ” A plate if filled to the rim with sweet pastries and desserts. ” “A cat laying next to a pair of shoes on a hard wood floor. ”
@@
Figure 8: The images below the corresponding MS-COCO captions are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right) under w4a8 with Stable-Diffusion-v1-4. Different key details are highlighted using rectangles.
Text-guided generation. In this experiment, we generate high-resolution images of $512\times 512$ pixels using Stable Diffusion-v1-4 with 50 steps PLMS sampler [51], and SD-XL-turbo with only a single step Euler sampler [31]. Moreover, higher resolution images of $1024\times 1024$ are also generated from SD-XL-base-1.0 with 50 steps Euler sampler. cfg is fixed to the default 7.5 for Stable Diffusion and SD-XL as the trade-off between generation quality and diversity. However, we do not use classifier-free guidance [23] to save memory, the same as the original paper [78]. In Tab. V, compared to Q-Diffusion, our approach achieves an FID reduction of 1.20, 5.61, and 5.50 under the w4a8 configuration for Stable Diffusion, SD-XL, and SD-XL-turbo, respectively. Notably, our FID on w8a8 and sFID on w4a32 are even lower than those of the full precision model. The table also shows our significantly lower quantization error than Q-Diffusion. For example, DS obtains an 8.16 SQNR upswing for w4a8 Stable Diffusion. Benefiting from this, the visual effect of our generated images  [^12] More qualitative comparisons can be found in Sec. H. is closer to the FP model with better quality than Q-Diffusion (examples in Fig. 8). All of these clues strongly reveal the value of maintaining temporal features. More experiments for few-step distilled model FLUX.1-Schnell [42] can also be found in Sec. F.
5.3 Efficiency Discussion
In this section, we study the inference efficiency of our quantized diffusion models. In the following, we discuss all components in our quantization framework in the same order as Sec. 4.4.
Specifically, we focus on popular T2I models, e.g., Stable Diffusion-v1-4 [74], SD-XL-base-1.0 [71] with w8a8 quantization and deploy their quantized counterparts on Intel® Xeon® Gold 6448Y Processor with the OpenVino [26] framework. To demonstrate the acceleration on GPU, we also evaluate the latency on Nvidia H800 GPU with quantized convolutions and multiplications harnessing CUTLASS [33] with our customized kernels. An overview can be seen from Fig. 10 and Fig. 9. In Sec. G, we additionally deploy Stable-Diffusion on the edge and mobile devices, e.g., Nvidia Jetson Orin Nano and iPhone 15 Pro Max.
Figure 9: Memory cost ablation for the UNet in Stable Diffusion and SD-XL.
@@ 5.3 Efficiency Discussion
Figure 10: Latency ablation for Stable Diffusion and SD-XL. BRECQ [45] is selected as our baseline. We set the batch size to 1 here. Full Prec. is implemented in w32a32 OpenVino and Pytorch.
5.3.1 Efficiency of TM

### v4

@@ IV-A Disturbance Observations
Figure 2: FID [21] on LSUN-Bedrooms $256\times 256$ [94] for LDM-4, with Gaussian noise applied to its temporal/non-temporal features.
Temporal feature disturbance. Having established the sensitivity of the temporal feature to disturbance, we thoroughly analyze its variations before and after the quantization of embedding layers and time embed in the Stable Diffusion model. Prior to this analysis, we introduce the temporal/non-temporal feature error at $t$ as defined by:
@@
Figure 3: (Left) Temporal feature disturbance. The green inflection points highlight the significant phenomenon of temporal feature disturbance. (Right) Temporal information mismatch ($i=11$). The coordinates of the inflection points on the blue curve can be denoted as $(t,t+\delta_{t,i})$. It indicates $\widehat{\mathbf{X}_{t,i}^{temporal}}$ exhibits the highest similarity with $\mathbf{X}_{t+\delta_{t,i},i}^{temporal}$. To be noted, for any point on the polylines in the above figures, assuming its x-axis is $t_{0}\in[1,\ldots,T]$, we pursue $T-t_{0}+1$ denoising processes to obtain its y-axis from the same random noise. Specifically, the model is set to full precision from $T$ to $t_{0}+1$ and only quantized precision at $t_{0}$.
@@ IV-B Disturbance Impacts
Temporal information mismatch. Obviously, temporal feature disturbance alters the original embedded temporal information. Specifically, $\mathbf{X}_{t,i}^{temporal}$ is intended to correspond to timestep $t$. However, due to significant errors, the quantized model’s $\widehat{\mathbf{X}_{t,i}^{temporal}}$ is no longer accurately associated with $t$, resulting in what we term as temporal information mismatch:
$t\leftarrow\mathbf{X}_{t,i}^{temporal},\quad t\nleftarrow\widehat{\mathbf{X}_{t,i}^{temporal}}.$ (10)
Furthermore, as depicted in Fig. 3 (Right), we even observe a pronounced temporal information mismatch in Stable Diffusion. Specifically, the temporal feature generated by the quantized model at timestep $t$ exhibits a divergence from that of the full-precision model at the corresponding timestep. Instead, it tends to align more closely with the temporal feature corresponding to $t+\delta_{t}$, importing wrong temporal information from $t+\delta_{t}$.
@@
where we apply disrupted temporal features to the model. Evidently, the deviation in the denoising trajectory intensifies as the number of denoising iterations increases, resulting in the final generated image failing to align accurately with $\mathbf{x}_{0}$. This evolution is illustrated in Fig. 4, where we maintain UNet in full precision, except for embedding layers and time embed. We sort out a detailed relationship between temporal feature errors and final generation quality in Sec. A.
@@
Figure 4: Denoising process of full-precision (Upper) and W4444A8888 quantized (Lower) Stable Diffusion $(T=50)$ under the same experiment settings, $\mathbf{x}_{0}$. It is noteworthy that, in the quantized model employed here, to showcase the impact of temporal features, only the layers generating temporal features are quantized, and the components unrelated to the generation of temporal features are maintained in full precision.
@@ IV-C Inducement Analyses
TABLE I: FID, sFID, and SQNR on LSUN-Bedrooms $256\times 256$ [94] for LDM-4. “Prev” represents BRECQ. Freeze denotes our trial. We conduct the same experiments with more baselines as “Prev” in Sec. B, which further validate our analysis.
@@
Unaware of finite activations within $h(\cdot)$ and $g_{i}(\cdot)$. We observe that, given $T$ as a finite positive integer, the set of all possible activation values for embedding layers and time embed is finite and strictly dependent on timesteps. Within this set, activations corresponding to the same layer display notable range variations across different timesteps  [^4] The details of range variations can be found in Sec. B of [24]. Previous methods [80, 84] mainly focus on finding the optimal calibration method for $\widehat{\mathbf{x}}_{t}$-related network components. Moreover, akin to the first inducement, their calibration is directly towards the Residual Bottleneck Block, which proves suboptimal  [^5] The evidence can be found in Sec. C of [24]. However, based on the finite activations, we can employ calibration methods, especially for these time-related activations, to better adapt to their range variations.
IV-D Quantization Framework
To mitigate temporal feature disturbance, we propose two different temporal feature maintenance strategies categorized into TIB-based and cache-based ones in Sec. IV-D1 and Sec. IV-D2, respectively. Finally, we incorporate both scenarios together with our selection strategy to further solve the problem in Sec. IV-D3. Our methods are outlined in Fig. 5.
IV-D1 TIB-based Maintenance
@@
$\mathop{\arg\,\min}\limits_{\mathcal{S}^{\mathbf{W}}_{g\circ h},\mathcal{Z}^{\mathbf{W}}_{g\circ h}}\sum_{i=1}^{n}\underbrace{\|g_{i}(h(t))-\widehat{g}_{i}(\widehat{h}(t))\|_{F}^{2}}_{\mathcal{L}^{temporal}_{t,i}},$ (12)
@@
Finite set calibration. To address the challenge posed by the wide span of activations within a finite set for the second inducement, we propose Finite Set Calibration (FSC) for activation quantization. This strategy employs $T$ sets of quantization parameters for activation, such as $\{(s_{T},z_{T}),\ldots,(s_{1},z_{1})\}$ for an arbitrary activation $\mathbf{X}$ within embedding layers and time embed. At timestep $t$, the quantization function for the $\mathbf{X}$ can be expressed as:
@@
where $s_{t}$ and $z_{t}$ are the corresponding scale and zero offset. Equipped with FSC, we can obtain $\mathcal{S}^{\mathbf{X}}_{g\circ h}$ and $\mathcal{Z}^{\mathbf{X}}_{g\circ h}$, which refer to all the quantization scales and zero offsets for the activations  [^7] We consider the outputs and inputs of every linear operator like [78, 19]. in $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$, respectively. To be noted, the calibration target for these activations is also aligned with the output of the TIB. Besides that, we find that Min-max [61] can achieve satisfactory results with high efficiency (more evidence in Sec. V-D2) for range estimation. In Sec. V-C1, we perform a detailed analysis to demonstrate that our method incurs a negligible extra cost for inference.
IV-D2 Cache-based Maintenance
@@
where $s_{t,i}$ and $z_{t,i}$ denote the quantization scale and zero offset for the pre-gained $\mathbf{X}_{t,i}^{temporal}$ (i.e., the output of $g_{i}\circ h$ at $t$). Finally, the obtained quantization parameters can be defined as $\mathcal{S}^{temporal}=\{s_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$ and $\mathcal{Z}^{temporal}=\{z_{t,i}\}_{t=1,\ldots,T\&i=1,\ldots,n}$. Specifically, in this strategy, we employ LSQ [11] to optimize every objective  [^8] From Eq. (14), it is obvious that we have to optimize $n\times T$ objectives across different indexes of temporal features. separately, which can be done in just a few minutes. During inference, this strategy enables us to get the temporal feature by reloading the cached items. In addition, we find that this approach can help improve latency in cases. More detailed analysis can be found in Sec. V-C2.
@@
IV-D3 Disturbance-aware Selection
Figure 6: Comparison of the temporal feature disturbance between TIB-based and Cache-based Maintenance with W4444A8888 quantized Stable Diffusion ($T=50$).
@@ IV-D3 Disturbance-aware Selection
Figure 7: Samples from W4444A8888 quantized Stable Diffusion on MS-COCO [49] caption with TIB-based Maintenance (Left), Cache-based Maintenance (Middle), and Disturbance-aware Selection (Right), where green rectangles present finer points than red rectangles. Numbers in the last row denote FID on the MS-COCO caption of the corresponding methods.
@@
V Experiments
We organize the experiments as follows: In Sec. V-A, we first demonstrate implementation details. Then, we exhibit the outstanding performance gained by our framework in Sec. V-B. Further, we also deploy our quantized models on various hardware to investigate their inference efficiency in Sec. V-C. Last, we conduct comprehensive ablative studies for the proposed framework in Sec. V-D. It is worth noting that we also discuss and conduct experiments for video generation in Sec. C. A guideline on when to use which proposed strategies and sub-$4$-bit quantization results are included in Sec. D and Sec. E, respectively.
V-A Implementation Details
Models and datasets. In this section, we conduct image generation experiments to evaluate the proposed quantization framework on various diffusion models: pixel-space diffusion model DDPM [22] for unconditional image generation, latent-space diffusion model LDM [73] for unconditional image generation and class-conditional image generation. We also apply our work to Stable Diffusion-v1-4 [73], SD-XL-base-1.0 [70], and SD-XL-turbo [77] for text-guided image generation. In our experiments, We use seven standard benchmarks: CIFAR-10 $32\times 32$ [42], LSUN-Bedrooms $256\times 256$ [94], LSUN-Churches $256\times 256$ [94], CelebA-HQ $256\times 256$ [32], ImageNet $256\times 256$ [9], FFHQ $256\times 256$ [34], and MS-COCO [49].
@@ V-A Implementation Details
Evaluation metrics. For each experiment, we evaluate the performance of diffusion models with Fréchet Inception Distance (FID) [21], and Signal-to-quantization-noise Ratio (SQNR) [66], which measures the quantization distortion in detail. In the case of LDM and text-guided generation experiments, we also include sFID [75], which better captures spatial relationships than FID. For ImageNet and CIFAR-10 experiments, we additionally provide Inception Score (IS) [75] as a reference metric. Further, in the context of text-guided experiments, we extend our evaluation to include the compatibility of image-caption pairs, employing the CLIP score [20]. The ViT-B/32 [10] is used as the backbone when computing the CLIP score. To ensure consistency in the reported outcomes, all results are derived from our implementation or from other papers, where experiments are conducted under conditions consistent with ours. More specifically, in the evaluation process of each experiment, we sample $50K$ images from DDPM or LDM, $30K$ images from Stable Diffusion or SD-XL-turbo, and $10K$ images from SD-XL. All experiments are conducted utilizing one $H800$ GPU and implemented with the PyTorch framework [67] without special claims.
V-B Performance Comparison
@@
TABLE IV: Quantization results for class-conditional image generation with LDM-4 on ImageNet $256\times 256$.
@@
TABLE V: Quantization results for text-guided image generation with Stable Diffusion-v1-4 and SD-XL on MS-COCO captions.
@@
Unconditional generation. In the experiments conducted on the LDM with DDIM sampler [81], we maintain the same experimental settings as presented in  [73], including the number of steps, variance schedule, and classifier-free guidance scale (denoted by eta and cfg in the following, respectively). As shown in Tab. II, the FID performance differences relative to the full precision (FP) model are all within $0.7$ for all settings employing DS. Specifically, on the CelebA-HQ $256\times 256$ dataset, DS exhibits an FID reduction of $\mathbf{6.74}$, an sFID reduction of $\mathbf{6.79}$, and an SQNR increase of $\mathbf{4.09}$ in the W4444A8888 setting compared to the current algorithms. It is noticeable that existing methods, whether in $4$-bit or $8$-bit, show significant performance degradation when compared to the FP model on face datasets like CelebA-HQ $256\times 256$ and FFHQ $256\times 256$, whereas our TM/CM/DS shows almost no performance degradation. Importantly, DS also achieves significant performance improvement on the LSUN-Bedrooms $256\times 256$ in the W4444A8888 setting, with FID and sFID reductions of $\mathbf{2.33}$ and $\mathbf{7.67}$ compared to PTQD [19], respectively. Regarding LDM-8 on LSUN-Churches $256\times 256$, we attribute the moderate improvement, compared to other datasets. We believe that the use of the LDM-8 model with a downsampling factor of $8$ may be more quantization-friendly. Existing methods have already achieved satisfactory results on this dataset. Nonetheless, our method still approaches the performance of the FP model more closely than existing methods.
Besides the experiments on LDM, we have also conducted experiments with DDPM on CIFAR-10 $32\times 32$. As shown in Tab. IV, our methods still achieve comprehensive improvements in terms of IS, FID, and SQNR compared to the existing advanced methods. However, due to the lower resolution and simplicity of the images in this dataset, existing methods show minimal performance degradation, so the results we obtain may not be as pronounced.
Class-conditional generation. On the ImageNet $256\times 256$ dataset, we employed a denoising process with a $20$-step DDIM sampler [81], setting eta and cfg to $0.0$ and $3.0$, respectively. In Tab. IV, compared to PTQD, our method achieves an FID reduction of $\mathbf{2.09}$ on W8888A8888, and a $\mathbf{5.35}$ sFID decrease on W4444A8888. Under all the conditions with DS, we observe an improvement of over $\mathbf{9}$ in IS. Particularly noteworthy is that, across various quantization settings, our method consistently achieved lower FID and higher SQNR compared to the FP model.
@@
“Elephant displayed in window of city building near roadway. ” “A plate if filled to the rim with sweet pastries and desserts. ” “A cat laying next to a pair of shoes on a hard wood floor. ”
@@
Figure 8: The images below the corresponding MS-COCO captions are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right) under W4444A8888 with Stable Diffusion-v1-4. Different key details are highlighted using rectangles.
@@
Figure 9: Latency for Stable Diffusion and SD-XL. BRECQ [46] is selected as our baseline. We set the batch size to $1$ here. Full Prec. is implemented in W32323232A32323232 OpenVino and Pytorch.
Figure 10: Memory costs for the UNet in Stable Diffusion and SD-XL.
Text-guided generation. In this experiment, we generate high-resolution images of $512\times 512$ pixels using Stable Diffusion-v1-4 with $50$ steps PLMS sampler [51], and SD-XL-turbo with only a single step Euler sampler [33]. Moreover, higher resolution images of $1024\times 1024$ are also generated from SD-XL-base-1.0 with a $50$-step Euler sampler. cfg is fixed to the default $7.5$ for Stable Diffusion and SD-XL as the trade-off between generation quality and diversity. However, we do not use classifier-free guidance [23] to save memory, the same as the original paper [77]. In Tab. V, compared to Q-Diffusion, our approach achieves an FID reduction of $\mathbf{1.20}$, $\mathbf{5.61}$, and $\mathbf{5.50}$ under the W4444A8888 configuration for Stable Diffusion, SD-XL, and SD-XL-turbo, respectively. Notably, our FID on W8888A8888 and sFID on W4444A32323232 are even lower than those of the full precision model. The table also shows our significantly lower quantization error than Q-Diffusion. For example, DS obtains an $\mathbf{8.16}$ SQNR upswing for W4444A8888 Stable Diffusion. Benefiting from this, the visual effect of our generated images is closer to the FP model with better quality than Q-Diffusion (examples in Fig. 8 and Sec. H). All of these clues strongly reveal the value of maintaining temporal features. More experiments for few-step FLUX.1-Schnell [43] can also be found in Sec. F.
V-C Efficiency Discussion
In this section, we study the inference efficiency of our quantized diffusion models. We discuss all components in our framework in the same order as Sec. IV-D.
Specifically, we focus on popular T2I models, e.g., Stable Diffusion-v1-4 [73], SD-XL-base-1.0 [70] with W8888A8888 quantization, and deploy their quantized counterparts on Intel® Xeon® Gold $6448Y$ Processor with the OpenVino [28] framework. To demonstrate the acceleration on GPU, we also evaluate the latency on NVIDIA $H800$ GPU with quantized convolutions and multiplications, harnessing CUTLASS [35] with our customized kernels. An overview can be seen from Fig. 9 and Fig. 10. In Sec. G, we additionally deploy Stable Diffusion on the edge and mobile devices, e.g., NVIDIA Jetson Orin Nano and iPhone $15$ Pro Max.
V-C1 Efficiency of TM